OpenStudy (anonymous):

What is the slope of the line identified by 6y = -3(x - 4)?

OpenStudy (solomonzelman):

divide both sides by 6. 6y = -3(x - 4) ----> y = -1/2x - 2/3 y=mx+b m is the slope, compare... m= -1/2,